# Central Path Proximal Policy Optimization

### v1

@@ Abstract
In constrained Markov decision processes, enforcing constraints during training is often thought of as decreasing the final return. Recently, it was shown that constraints can be incorporated directly in the policy geometry, yielding an optimization trajectory close to the central path of a barrier method, which does not compromise final return. Building on this idea, we introduce Central Path Proximal Policy Optimization (C3PO), a simple modification of the PPO loss that produces policy iterates, which stay close to the central path of the constrained optimization problem. Compared to existing on-policy methods, C3PO delivers improved performance with tighter constraint enforcement, suggesting that central path-guided updates offer a promising direction for constrained policy optimization.
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated impressive capabilities across a wide range of domains, yet real-world applications increasingly demand more than just reward maximization. In many real-world high-stakes environments—such as autonomous vehicles, healthcare, or robotic manipulation—agents must also avoid violating domain-specific safety or resource constraints. This motivates the study of constrained Markov decision processes (CMDPs), an extension of the standard RL framework that imposes expected cost constraints alongside the goal of reward maximization Altman (1999). By treating feasibility and reward objectives separately, CMDPs provide a principled framework for specifying agent behavior in complex environments.
CMDPs are especially relevant in deep reinforcement learning, where designing reward functions that reliably encode which states to avoid is notoriously difficult. Prior work has emphasized the importance of explicit constraint modeling for safe behavior. For instance, Ray et al. (2019) argue that CMDPs offer a natural formalism for benchmarking safe exploration in deep RL and introduce the Safety Gym suite to evaluate algorithms based on both task performance and cumulative safety cost. Similarly, Roy et al. (2022) demonstrate that task specifications expressed via constraints are often more natural and easier to design, especially as tasks increase in complexity, e.g. in finetuning LLMs Dai et al. (2023).
@@
In this context, there is a growing need for simple, scalable, and effective algorithms for solving CMDPs—ideally with properties similar to widely used algorithms such as proximal policy optimization (PPO; Schulman et al. (2017b)). PPO’s robustness, ease of implementation, and scalability have made it the method of choice in many deep RL and RLHF pipelines Ouyang et al. (2022). We aim to extend these strengths to the constrained setting by developing an algorithm that shares PPO’s practical benefits while enforcing constraints in a principled CMDP framework. Specifically, we seek to achieve high final reward while approximately satisfying constraints, at least at convergence.
To frame this problem, we distinguish between two commonly conflated settings in constrained RL: (i) safe exploration, where constraints must be satisfied throughout training, and (ii) safe convergence, where only the final policy is required to satisfy the constraints. Much of the literature has focused on the former, motivated by safety-critical applications in the real world. In contrast, safe convergence—where exploration may be unsafe—better reflects settings like simulation-based training or alignment finetuning Dai et al. (2023).Typically, ensuring safety during training is considered to decrease the final performance achieved by an algorithm. We show the contrary and present an algorithm that exhibits strict feasibility during training as well as reliable feasibility and high return at convergence.
@@
In nonlinear CMDPs, the constraint surface is typically curved and nonconvex in policy parameter space. Converging prematurely or oscillating near the constraint boundary during training—as is often the case with Lagrangian or most trust region methods—can lead to unreliable constraint satisfaction at convergence. Furthermore, it can lead the optimizer to local optima that satisfy the constraints but fail to achieve high reward, see Figure 1. Penalty and barrier methods avoid this by maintaining a feasible trajectory toward the constraint surface, yielding feasible solutions more reliably. However, barrier methods introduce bias (Müller and Cayci, 2024), meaning the optimization problem obtained by adding a barrier penalty does not have the same solution set as the original problem, which can lead to degraded reward in policy optimization Milosevic et al. (2025). Barrier methods either require careful tuning or an interior point approach Liu et al. (2020) to avoid harming reward performance.
The recently proposed C-TRPO (Milosevic et al., 2025) addresses these challenges by combining the strengths of trust-region and barrier methods by deriving a barrier-inspired trust-region formulation using strictly feasible trust regions. This results in an algorithm that acts like a barrier method with an adaptively receding barrier. This introduces no bias in the optimal feasible solution but still produces strictly feasible policies just like a barrier method. Further, C-TRPO produces policies, which are close to the regularization path obtained by altering the regularization strength, which is commonly known as the central path Boyd and Vandenberghe (2004); Müller and Cayci (2024). This nicely illustrates how the constraints are incorporated in the algorithm’s geometry and ensures that C-TRPO and variants thereof produce policies which don’t prematurely approach the constraint surface.
However, C-TRPO’s scalability remains limited due to computational overhead and sensitivity to large network architectures or batch sizes introduced by the TRPO-inspired update, and the update is defined only in the feasible set. To address the need for a simple and scalable CMDP solver, we propose a proximal version of C-TRPO. It also follows the central path, which we therefore call Central Path Proximal Policy Optimization (C3PO). C3PO is a minibatch-based method that approximates the C-TRPO update using an exact penalty formulation, combining the simplicity and efficiency of PPO-style updates with the feasible geometry of central path methods. At its core, C3PO leverages the central path property of natural policy gradients to gradually guide the policy toward the constraint surface without inducing oscillations or premature convergence. The result is a practical algorithm that retains high reward performance while satisfying constraints at convergence, and which has the potential to scale well to large neural networks and modern deep RL settings.
@@ 2 Background
We consider the infinite-horizon discounted constrained Markov decision process (CMDP) and refer the reader to Altman (1999) for a general treatment. The CMDP is given by the tuple $\mathcal{M}\cup\mathcal{C}$, consisting of a finite MDP $\mathcal{M}$ and a set of constraints $\mathcal{C}$. The finite MDP =$\mathcal{M}=\{\mathcal{S},\mathcal{A},P,r,\mu,\gamma\}$ is defined by a finite state-space $\mathcal{S}$, a finite action-space $\mathcal{A}$, a transition kernel $P\colon\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$, an extrinsic reward function $r\colon\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, an initial state distribution $\mu\in\Delta_{\mathcal{S}}$, and a discount factor $\gamma\in[0,1)$. The space $\Delta_{\mathcal{S}}$ is the set of categorical distributions over $\mathcal{S}$. Further, $\mathcal{C}=\{(c_{i},b_{i})\}_{i=1}^{m}$ is a set of $m$ constraints, where $c_{i}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are the cost functions and $d_{i}\in\mathbb{R}$ are the cost thresholds. An agent interacts with the CMDP by selecting a policy $\pi\in\Pi$. Given $\pi$, the value function $V_{r}^{\pi}\colon\mathcal{S}\to\mathbb{R}$, action-value function $Q_{r}^{\pi}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, and advantage function $A_{r}^{\pi}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ associated with the reward function $r$ are defined as
$V_{r}^{\pi}(s)\coloneqq(1-\gamma)\,\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\Big{|}s_{0}=s\right],$
$Q_{r}^{\pi}(s,a)\coloneqq(1-\gamma)\,\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\Big{|}s_{0}=s,a_{0}=a\right],$
@@
The expectations are taken over trajectories of the Markov process, meaning with respect to the initial distribution $s_{0}\sim\mu$, the policy $a_{t}\sim\pi(\cdot|s_{t})$ and the state transition $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. They are defined analogously for the $i$-th cost and denoted as $V_{c_{i}}^{\pi}(s)$, $Q_{c_{i}}^{\pi}(s,a)$ and $A_{c_{i}}^{\pi}(s,a)$.
@@
Every stationary policy $\pi$ induces a discounted state-action occupancy measure $\rho_{\pi}\in\mathcal{K}\subset\Delta_{\mathcal{S}\times\mathcal{A}}$, which indicates the relative frequencies of visiting a state-action pair, discounted by how far the event lies in the future. We will refer to this measure as the state-action occupancy for short.
Definition 2.1.
The state-action occupancy $\rho_{\pi}$ is defined as
$\rho_{\pi}(s,a)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s)\pi(a|s),$ (2)
where $\mathbb{P}_{\pi}(s_{t}=s)$ is the probability of observing the environment in state $s$ at time $t$ given $\pi$.
Note that similar measures can be introduced for the average-reward setting Zahavy et al. (2021), and for continuous state-action spaces Laroche and Des Combes (2023). For any state-action measure $\rho$ we obtain the associated policy via conditioning, meaning
$\pi(a|s)\coloneqq\frac{\rho(s,a)}{\sum_{a^{\prime}}\rho(s,a^{\prime})}.$ (3)
Assumption 2.2 (State Exploration).
For any policy $\pi\in\operatorname{int}(\Delta_{\mathcal{A}}^{\mathcal{S}})$ we have $\rho_{\pi}(s)>0$ for all $s\in\mathcal{S}$.
Under 2.2, this provides a one-to-one correspondence between the space of policies $\Pi$ and the space of state-action occupancies $\mathcal{K}$, which forms a convex polytope inside of $\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ Kallenberg (1994); Mei et al. (2020). These properties justify the linear programming (LP) approach to solving finite CMDPs Altman (1999), where problem 1 is reformulated as
@@
which can be solved to obtain an optimal occupancy $\rho^{*}$ using LP solution methods. The optimal policy can be extracted using relation 3.
In the function approximation setting, approach 4 is not applicable, which has prompted a large body of research in on-policy policy optimization methods. In the analysis of on-policy methods–including for standard MDPs–the policy advantage plays an important role. The policy (reward) advantage and policy cost advantage are defined as
@@
It features prominently in policy optimization methods, as it approximates the performance difference between two nearby policies, i.e. $\mathbb{A}_{r}^{\pi_{k}}(\pi)\approx R(\pi)-R(\pi_{k})$ if $\pi_{k}\approx\pi$, and for the policy cost advantage analogously.
@@ 2.1 Policy Optimization Methods for Constrained Reinforcement Learning
Constrained policy optimization (CPO) is a modification of trust region policy optimization (TRPO; Schulman et al. (2017a)), where the classic trust region is intersected with the set of safe policies Achiam et al. (2017). At each iteration $k$, the policy of the next iteration $\pi_{k+1}$ is obtained through the solution of
@@
The appeal of this reformulation is that one can obtain an unconstrained problem that gives the same solution set for $\lambda$ chosen large enough Zhang et al. (2022) and by employing a PPO-like loss. A similar approach is taken by Dey et al. (2024), where $\max\{0,\cdot\}$ is replaced by a softplus.
@@
Where equation 8 incorporates the safety by intersecting the trust region with the set of safe policies, an alternative approach was taken in Milosevic et al. (2025) where the geometry was modified such that the resulting trust region automatically consists of safe policies. To this end, C-TRPO proceeds as TRPO but with the usual divergence augmented by a barrier term, meaning
@@
as the barrier divergence [^1] Note the similarity to the unbiased KL-Divergence estimator $\hat{D}_{KL}=\frac{\pi(a|s)}{\pi_{k}(a|s)}-\log\frac{\pi(a|s)}{\pi_{k}(a|s)}-1$, $\beta$ is a positive safety parameter, and $b=d-C(\pi_{k})$ is its cost budget.
@@ 3 Central Path Proximal Policy Optimization
Figure 2: The working principle behind C3PO’s exact penalty approach: As the iterate moves closer towards the constraint (A-C), C3PO’s ReLU-penalty pulls away at a slower rate $10<w<$ 1, mimicking C-TRPO’s barrier divergence. This rate is defined as a function of $\delta_{\textup{B}}$ (D), see main text. While C-TRPO’s barrier penalty is undefined if either $d\leq C(\pi_{k})$ or $d\leq C(\pi_{k+1})$, C3PO’s ReLU-penalty is defined everywhere.
@@
Note that the problem Equation 10 is concave-convex in $\pi$. Hence, this is a special case of the general exactness result Theorem B.1. ∎
@@
The update Equation 11 is still undefined outside the feasible set of the barrier divergence constraint. Since we use the barrier divergence only to define the feasible solution set of the update, we can replace it with another function, as long as it defines the same feasible set. More precisely, this can be achieved with an equivalent linear constraint that is zero where $D_{\textup{B}}(\pi,\pi_{k})=\delta_{\textup{B}}$ for positive cost advantages. The C3PO algorithm approximates update 11 as
@@
For $0\leq\mathbb{A}_{c}^{\pi_{k}}(\pi)<d-C(\pi_{k})$, there exist $w$ and $\delta_{\textup{B}}$ for which the solution sets of problems 10, 11 and 12 agree.
The new update expresses the same constraint using a linear ReLU-penalty. The rate $w$ is a new hyper-parameter and we refer to Appendix B for a proof of Proposition 3.2. Since the original problem’s penalty function is not defined outside the interior of the feasible set, we must handle the case $C(\pi_{k})\geq d$ differently, which is taken care of by the $\min(b,\cdot)$ term: For $b<0$, problem 12 reduces to the P3O Zhang et al. (2022) objective Eq. 7. Finally, the additional KL-constraint is approximated as in PPO Schulman et al. (2017b). The resulting loss only consists of the PPO loss and an additional loss term which is a function of the policy cost advantage estimate. Let $r(\theta)=\frac{\pi_{\theta}(a|s)}{\pi_{k}(a|s)}$ denote the likelihood ratio of the optimized and last behavior policies and let
@@
The penalty coefficient remains a hyperparameter, which can be flexibly scheduled to solve either setting (i) or (ii), as we demonstrate in Section 4, where we use a linear schedule to achieve high final performance across multiple tasks. The final method is summarized in Algorithm 1.
@@
4:     Estimate reward advantage $\hat{A}^{r}_{t}$ and cost advantages $\hat{A}^{c_{i}}_{t}$ using GAE-$\lambda$ Schulman et al. (2018)
@@
C3PO is a superset of P3O Zhang et al. (2020). More precisely, if we set $w=1$ in C3PO, we obtain the P3O loss exactly. Further, C3PO is conceptually similar to P2BPO Dey et al. (2024), in using a more conservative version of the P3O loss, but C3PO does not use a penalty with a fixed location at the constraint, but a moving penalty which recedes as the iterate gets closer to the constraint. This allows C3PO to approach the optimal feasible solution without regularization bias.

### v2

@@ Abstract
In constrained Markov decision processes, enforcing constraints during training is often thought of as decreasing the final return. Recently, it was shown that constraints can be incorporated directly into the policy geometry, yielding an optimization trajectory close to the central path of a barrier method, which does not compromise final return. Building on this idea, we introduce Central Path Proximal Policy Optimization (C3PO), a simple modification of the PPO loss that produces policy iterates, that stay close to the central path of the constrained optimization problem. Compared to existing on-policy methods, C3PO delivers improved performance with tighter constraint enforcement, suggesting that central path-guided updates offer a promising direction for constrained policy optimization.
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated impressive capabilities across a wide range of domains, yet real-world applications increasingly demand more than just reward maximization. In many real-world high-stakes environments agents must also avoid violating domain-specific safety or resource constraints. This motivates the study of constrained Markov decision processes (CMDPs), an extension of the standard RL framework that imposes expected cost constraints alongside the goal of reward maximization (Altman, 1999). By treating feasibility and reward objectives separately, CMDPs provide a principled framework for specifying agent behavior in complex environments.
CMDPs are especially relevant in deep reinforcement learning settings, where the design of reward functions that lead to safe behavior is difficult. Prior work has emphasized the importance of explicit constraint modeling in reinforcement learning for safe exploration (Ray et al., 2019) and complex task specification (Roy et al., 2022), where constraints can be more natural and easier to design, including finetuning LLMs for harmlessness (Dai et al., 2024).
@@
In this context, there is a growing need for simple, scalable, and effective algorithms for solving CMDPs, ideally with properties similar to widely used algorithms such as proximal policy optimization (PPO; Schulman et al. (2017b)). PPO’s robustness, ease of implementation, and scalability have made it the method of choice in many deep RL and RLHF pipelines (Ouyang et al., 2022). We aim to extend these strengths to the constrained setting by developing an algorithm that shares PPO’s practical benefits while enforcing constraints in a principled CMDP framework. Specifically, we seek to achieve high final reward while approximately satisfying constraints, at least at convergence.
To frame this problem, we distinguish between two commonly conflated settings in constrained RL: (i) safe exploration, where constraints must be satisfied throughout training, and (ii) safe convergence, where only the final policy is required to satisfy the constraints. Much of the literature has focused on the former, motivated by safety-critical applications in the real world. The dominant approach in this setting is model-based safe RL, which can provide strong safety and stability guarantees (Berkenkamp et al., 2017; As et al., 2025). In contrast, safe convergence reflects settings like simulation-based training or alignment finetuning (Dai et al., 2024). Typically, ensuring safety during training is considered to decrease the final performance achieved by an algorithm. We show the contrary and present an algorithm that exhibits strict feasibility during training as well as reliable feasibility and high return at convergence.
@@
In nonlinear CMDPs, the constraint surface is typically curved and nonconvex in policy space. Converging prematurely or oscillating near the constraint boundary during training can lead to unreliable constraint satisfaction at convergence. Furthermore, it can lead the iterates to local optima that satisfy the constraints but fail to achieve high reward, see Figure 1. Penalty and barrier methods address this problem by maintaining a feasible trajectory toward the constraint surface, yielding feasible solutions more reliably. However, barrier methods introduce bias (Müller and Cayci, 2024), meaning the optimization problem obtained by adding a barrier penalty does not have the same solution set as the original problem, which can lead to degraded reward in policy optimization (Milosevic et al., 2025). Barrier methods either require careful tuning or an interior point approach (Liu et al., 2020) to avoid harming reward performance.
The recently proposed C-TRPO (Milosevic et al., 2025) addresses these challenges by combining the strengths of trust-region and barrier methods by deriving a barrier-inspired trust-region formulation using strictly feasible trust regions. This results in an algorithm that acts like a barrier method with an adaptively receding barrier, introducing no regularization bias as a result. This is achieved by producing policies, which are close to the regularization path obtained by altering the regularization strength, which is commonly known as the central path (Boyd and Vandenberghe, 2004).
C-TRPO’s scalability is limited due to computational overhead introduced by the TRPO-inspired update, and the update is defined only in the feasible set. To address the need for a simple and scalable CMDP solver, we propose a proximal version of C-TRPO. It also follows the central path, and consequently we call it Central Path Proximal Policy Optimization (C3PO). C3PO is a minibatch-based method that approximates the C-TRPO update using an exact penalty formulation, combining the simplicity and efficiency of PPO-style updates with the feasible geometry of central path methods. At its core, C3PO leverages the central path property of natural policy gradients to gradually guide the policy toward the constraint surface without inducing oscillations or premature convergence.
@@ 2 Background
We consider the infinite-horizon discounted constrained Markov decision process (CMDP) and refer the reader to Altman (1999) for a general treatment. The CMDP is given by the tuple $\mathcal{M}\cup\mathcal{C}$, consisting of a finite MDP $\mathcal{M}$ and a set of constraints $\mathcal{C}$. The finite MDP $\mathcal{M}=\{\mathcal{S},\mathcal{A},P,r,\mu,\gamma\}$ is defined by a finite state-space $\mathcal{S}$, a finite action-space $\mathcal{A}$, a transition kernel $P\colon\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$, an extrinsic reward function $r\colon\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, an initial state distribution $\mu\in\Delta_{\mathcal{S}}$, and a discount factor $\gamma\in[0,1)$. The space $\Delta_{\mathcal{S}}$ is the set of categorical distributions over $\mathcal{S}$. Further, $\mathcal{C}=\{(c_{i},b_{i})\}_{i=1}^{m}$ is a set of $m$ constraints, where $c_{i}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are the cost functions and $d_{i}\in\mathbb{R}$ are the cost thresholds. An agent interacts with the CMDP by selecting a policy $\pi\in\Pi$ and collecting trajectories $\tau=(s_{0},a_{0},...s_{T})$. Let $R(\tau)=\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$. Given $\pi$, the value function, action-value function, and advantage function associated with the reward function $r$ are defined respectively as
$V_{r}^{\pi}(s)\coloneqq(1-\gamma)\,\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\Big{|}s_{0}=s\right],\quad Q_{r}^{\pi}(s,a)\coloneqq(1-\gamma)\,\mathbb{E}_{\tau\sim\pi}\left[R(\tau)\Big{|}s_{0}=s,a_{0}=a\right],$
@@
The expectations are taken over trajectories of the Markov process, meaning with respect to the initial distribution $s_{0}\sim\mu$, the policy $a_{t}\sim\pi(\cdot|s_{t})$ and the state transition $s_{t+1}\sim P(\cdot|s_{t},a_{t})$. $V_{c_{i}}^{\pi}(s)$, $Q_{c_{i}}^{\pi}(s,a)$ and $A_{c_{i}}^{\pi}(s,a)$ are defined analogously for the $i$-th cost $c_{i}$ instead of $r$.
@@
Every stationary policy $\pi$ induces the occupancy measures $\rho_{\pi}(s)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}_{\pi}(s_{t}=s)$, and $\rho_{\pi}(s,a)\coloneqq\rho_{\pi}(s)\pi(a|s)$ which indicate the relative frequencies of visiting a state(-action) pair, discounted by how far the event lies in the future. The classical, linear programming (LP) approach to solving finite CMDPs (Altman, 1999), reformulates problem 1 as
@@
which can be solved using LP solution methods to obtain an optimal occupancy measure $\rho^{*}$. Here, $\mathcal{K}$ is a set of linear constraints that $\rho_{\pi}$ must satisfy (Kallenberg, 1994; Mei et al., 2020), sometimes referred to as the Bellman flow equations. Finally, an optimal policy can be extracted by conditioning $\pi^{*}(a|s)=\rho^{*}(s,a)/\sum_{a^{\prime}}\rho^{*}(s,a^{\prime})$.
In the function approximation setting, approach 2 is not applicable, which has prompted a large body of research in on-policy policy optimization methods. However, it can be leveraged to derive general constrained RL algorithms (Milosevic et al., 2025). In the analysis of on-policy methods (including for standard MDPs) the policy advantage plays an important role. In CMDPs, the policy advantages are defined as
@@
They play an important role in policy optimization, as they approximate the performance difference between two nearby policies with respect to the reward $\mathbb{A}_{r}^{\pi_{k}}(\pi)\approx R(\pi)-R(\pi_{k})$ if $\pi_{k}\approx\pi$ and analogously for the cost function.
@@ 2.1 Policy Optimization Methods for Constrained Reinforcement Learning
Constrained policy optimization (CPO) is a modification of trust region policy optimization (TRPO; Schulman et al. (2017a)), where the classic trust region is intersected with the set of safe policies (Achiam et al., 2017). At each iteration $k$, the policy of the next iteration $\pi_{k+1}$ is obtained through the solution of
@@
The appeal of this reformulation is that one can obtain an unconstrained problem that gives the same solution set for $\lambda$ chosen large enough (Zhang et al., 2022) and by employing a PPO-like loss. A similar approach was taken by Dey et al. (2024), where $\max\{0,\cdot\}$ was replaced with a softplus.
@@
Where equation 4 incorporates constraints by intersecting the trust region with the set of safe policies, an alternative approach was taken by Milosevic et al. (2025) where the geometry was modified such that the resulting trust region automatically consists of safe policies. To this end, C-TRPO proceeds as TRPO but with the usual divergence augmented by a barrier term, meaning
@@
as the barrier divergence, $\beta$ is a positive safety parameter, and $b=d-C(\pi_{k})$ is the cost budget.
@@
Other methods
So far we have focused on model-free, direct policy optimization methods, specifically trust-region and PPO-penalty based ones. However, it is important to note that model-based approaches, e.g. Berkenkamp et al. (2017); As et al. (2025), are also attractive due to their stability and safety guarantees, but require learning a model, which is not always feasible. Lagrangian methods are a widely adopted approach, where the CMDP is formulated as a primal-dual optimization problem. Optimizing the dual variable with stochastic gradient descent is a popular baseline (Achiam et al., 2017; Ray et al., 2019; Chow et al., 2019; Stooke et al., 2020). However, a naively optimized dual variable may cause oscillations and overshoot, and analyzing more nuanced update strategies is a subject of current research interest (Sohrabi et al., 2024). More recently, log-barrier approaches have attracted considerable research interest (Usmanova et al., 2024; Zhang et al., 2024; Dey et al., 2024) due to their algorithmic simplicity through the fixed penalty, but also due to recent rigorous treatments, see e.g. Ni and Kamgarpour (2024). However, working with an explicit penalty produces suboptimal policies w.r.t the original constrained MDP. Fixed penalties introduce an additional error, which has to be controlled, see for example Geist et al. (2019); Müller and Cayci (2024) for theory, and Liu et al. (2020) for a practical example of regularization bias.
@@ 3 Central Path Proximal Policy Optimization
Figure 2: The working principle behind C3PO’s exact penalty approach: As the iterate moves closer towards the constraint (A-C), C3PO’s ReLU-penalty pulls away at a slower rate $0<w<1$, mimicking C-TRPO’s barrier divergence. This rate is defined as a function of $\delta_{\textup{B}}$ (D), see main text. While C-TRPO’s barrier penalty is undefined if either $d\leq C(\pi_{k})$ or $d\leq C(\pi_{k+1})$, C3PO’s ReLU-penalty is defined everywhere.
@@
Note that the problem Equation 8 is concave-convex in $\pi$. Hence, this is a special case of the general exactness result Theorem˜B.1. ∎
@@
The update Equation 9 is still undefined outside the feasible set of the barrier divergence constraint. Since we use the barrier divergence only to define the feasible solution set of the update, we can replace it with another function, as long as it defines the same feasible set. More precisely, this can be achieved with an equivalent linear constraint that is zero where $D_{\textup{B}}(\pi,\pi_{k})=\delta_{\textup{B}}$ for positive cost advantages. The C3PO algorithm approximates update 9 as
@@
For $0\leq\mathbb{A}_{c}^{\pi_{k}}(\pi)<d-C(\pi_{k})$, there exist $w$ and $\delta_{\textup{B}}$ for which the solution sets of problems 8, 9 and 10 agree.
The new update expresses the same constraint using a linear ReLU-penalty. The rate $w$ is a new hyper-parameter and we refer to Appendix B for a proof of Proposition˜3.2. Since the original problem’s penalty function is not defined outside the interior of the feasible set, we must handle the case $C(\pi_{k})\geq d$ differently, which is taken care of by the $\min(b,\cdot)$ term: For $b<0$, problem 10 reduces to the P3O (Zhang et al., 2022) objective Eq. 5. Finally, the additional KL-constraint is approximated as in PPO (Schulman et al., 2017b). The resulting loss only consists of the PPO loss and an additional loss term which is a function of the policy cost advantage estimate. Let $r(\theta)=\frac{\pi_{\theta}(a|s)}{\pi_{k}(a|s)}$ denote the likelihood ratio of the optimized and last behavior policies and let
@@
The penalty coefficient remains a hyperparameter, which can be flexibly scheduled to solve CMDPs in the safe convergence regime, as shown in Section 4, where we use a linear schedule to achieve high final performance across multiple tasks. The final method is summarized in Algorithm 1.
@@
4:  Estimate reward advantage $\hat{A}^{r}_{t}$ and cost advantages $\hat{A}^{c_{i}}_{t}$ using GAE-$\lambda$ (Schulman et al., 2016)
@@
C3PO is a superset of P3O (Zhang et al., 2020). More precisely, if we set $w=1$ in C3PO, we obtain the P3O loss exactly. Further, C3PO is conceptually similar to P2BPO (Dey et al., 2024), in using a more conservative version of the P3O loss, but C3PO does not use a penalty with a fixed location at the constraint, but a moving penalty which recedes as the iterate gets closer to the constraint. This allows C3PO to approach the optimal feasible solution without regularization bias.